Redudjdjd:

Click an item in the list or group of pictures at the bottom of the problem and, holding the button down, drag it into the correct position in the answer box. Release your mouse button when the item is place. If you change your mind, drag the item to the trashcan. Click the trashcan to clear all your answers. Using the transformation T: (x, y) (x + 2, y + 1), find the distance named. Find the distance BB'

Redudjdjd:

1 attachment
KyledaGreat:

To find the distance BB', we need to determine the coordinates of points B and B' and then use the distance formula. Using the transformation T: (x, y) -> (x + 2, y + 1), we can find the coordinates of point B': B' = T(B) = (Bx + 2, By + 1) = (3 + 2, 2 + 1) = (5, 3) Therefore, the coordinates of point B are (3, 2) and the coordinates of point B' are (5, 3). Using the distance formula: distance BB' = √[(xB - xB')^2 + (yB - yB')^2] = √[(3 - 5)^2 + (2 - 3)^2] = √[(-2)^2 + (-1)^2] = √[4 + 1] = √5 Therefore, the distance BB' is √5 or approximately 2.236 units.